OpenStudy (anonymous):

Solve the equation for y and then find the value for y for each given x value. 3y-6y=-3 explain how to

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

3y - 6y = -3y. So... \[-3y=-3\]

OpenStudy (anonymous):

and then its y=-3/-3

OpenStudy (anonymous):

Oh, forgot to say it also says x=-4

OpenStudy (anonymous):

What would the equation be?

OpenStudy (anonymous):

Okay sorry it says find the equation for y and then find the value of y for each given x-value 3y-6x=-3 x=-4 y= ?

OpenStudy (anonymous):

So if you're given x = -4, use substitution.

OpenStudy (anonymous):

Sorry, I plugged it in wrong.

OpenStudy (anonymous):

3y + 24 = -3

OpenStudy (anonymous):

So y would equal -9?

OpenStudy (anonymous):

You know how to solve for y, right? ( in that format?)

OpenStudy (anonymous):

Yes, y = -9

OpenStudy (anonymous):

So if you want, you can plug back in, y=-9 and x=-4 back into your equation to double check,

OpenStudy (anonymous):

Thank you! :D